OpenStudy (anonymous):

Evaluate m0 - n2 for m = 2 and n = -1. A: -1 B: 0 C: 1 Help Please <3

OpenStudy (alfie):

There is no other possible answer?

OpenStudy (anonymous):

No :/. Just That.

OpenStudy (anonymous):

M^0 - N^2

OpenStudy (alfie):

Oh... pretty different. 2^0 = 1 (-1)^2 = 1 1-1 = 0

OpenStudy (anonymous):

Oh Okay Thank You (: